# Convergence of the open WASEP stationary measure without Liggett’s condition

### v1

@@ 3.1. Convergence of One-point Distributions
We begin by reminding ourselves of the expression for $\hat{\pi}_{t;v,w,u,r}^{(N),c}(\cdot)$. As before, we set $q=\exp{\left(-\frac{2}{\sqrt{N}}\right)}$. We will deal with the expression for the rescaled, time-reversed process $\hat{\mathbb{Y}}_{t}$, and the corresponding one-point measures $\hat{\pi}_{t;v,w,u,r}^{(N),c}(\cdot)$.
(3.3) $\hat{\pi}_{t;v,w,u,r}^{(N),c}(y)$ $=\frac{1}{2N}\pi_{q^{t}}^{(N),c}\left(1-\frac{y}{2N}\right)$
@@
×$\times\frac{(q,-q^{v+w+t},q^{u+v},-q^{v+r},-q^{u+w},q^{w+r},-q^{u+r-t})_{\infty}|(q^{i\sqrt{N}\theta})_{\infty}|^{2}}{|(q^{v+\frac{t}{2}+i\frac{\sqrt{N}\theta}{2}},-q^{w+\frac{t}{2}+i\frac{\sqrt{N}\theta}{2}},q^{u-\frac{t}{2}+i\frac{\sqrt{N}\theta}{2}},-q^{r-\frac{t}{2}+i\frac{\sqrt{N}\theta}{2}})_{\infty}|^{2}}$
We will only prove a bound in the interval $y\in[0,2N]$. The way that this bound can be applied to prove the full Lemma 3.2 is by using the symmetry (2.5) to turn this bound into in the proof by first proving our desired bound for $\hat{\pi}^{(N),c}_{t;v,w,u,r}(y)$ on the interval $y\in[0,2N]$, and then noting that the desired bound on the interval $[2N,4N]$ is significantly weaker than the bound on the interval $[0,2N]$ which arises from $\hat{\pi}^{(N),c}_{t;w,v,r,u}(4N-y)$. Therefore, we first prove the appropriate bound on (3.3) in the interval $r\in[0,2N]$, and then use this identity to reach the remaining $r\in[2N,4N].$ We take the logarithm of the one-point measure,
(3.4) $\log{\left(\sqrt{1-\frac{y}{4N}}\hat{\pi}_{t;v,w,u,r}^{(N),c}(y)\right)}$ $=-\log{\left(4\pi N\sqrt{\frac{y}{N}}\right)}+\mathcal{A}(\sqrt{y})+E^{\mathcal{A}}_{N}(y)+E^{\mathcal{E}}_{N}(y)$
@@
$\mathcal{A}_{N}(\sqrt{y})$ $:=\mathcal{A}^{+}_{N}\left[1,u+v,w+r,i\frac{\sqrt{y}}{2},-i\frac{\sqrt{y}}{2}\right]$
$+\mathcal{A}^{-}_{N}\left[v+w+t,v+r,u+w,u+r-t\right]$
$-\mathcal{A}_{N}^{+}\left[u+v+w+r,v+i\frac{\sqrt{y}}{2}+\frac{t}{2},u+i\frac{\sqrt{y}}{2}-\frac{t}{2},v-i\frac{\sqrt{y}}{2}+\frac{t}{2},u-i\frac{\sqrt{y}}{2}-\frac{t}{2}\right]$
$-\mathcal{A}_{N}^{-}\left[w+i\frac{\sqrt{y}}{2}+\frac{t}{2},r+i\frac{\sqrt{y}}{2}-\frac{t}{2},w-i\frac{\sqrt{y}}{2}+\frac{t}{2},r-i\frac{\sqrt{y}}{2}-\frac{t}{2}\right].$
@@
$E_{N}^{\theta}(y)$ $:=\sqrt{N}\arccos{\left(1-\frac{y}{2N}\right)}-\sqrt{y}.$
are a concise way of containing the logarithm of the q-pochammer symbols in $\hat{\pi}_{t;v,w,u,r}^{(N),c}(y)$. The notation $\mathcal{A}^{+}$ and $\mathcal{A}^{-}$ was defined in (2.13), and $E_{N}^{\theta}(y)$ arises as a concise expression for the error terms in the Taylor expansion $\sqrt{N}\theta_{y}^{(N)}=\sqrt{y}+E_{N}^{\theta}(y)$.
We can simplify the first two terms on the right-hand side of (3.4) considerably by expanding $\mathcal{A}(\sqrt{y})$ through the definitions of $\mathcal{A}^{+}[z_{1},...,z_{k}]$ and $\mathcal{A}^{-}[z_{1},...,z_{k}]$ from (2.13), and then canceling where appropriate. The result of that calculation yields
@@
$\log\left(N^{u+v}\sqrt{1-\frac{y}{4N}}\hat{\pi}_{t;v,w,u,r}^{(N),c}(y)\right)$ $=\log{(\mathfrak{p}_{t}^{c}(y))}+\log{(C_{u,v,w,r})}+E_{N}^{\mathcal{A}}(y)+E_{N}^{\mathcal{E}}(y).$
@@
    For all $\eta>1$ there exists $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi>0$ such that for all $N>N_{0}$ and $y\in\hat{S}^{(N),c}=[0,2N]$,
@@
    $|\textsc{Error}_{1}^{\pm}[\kappa,z]|\leq C(\kappa+\kappa^{b})\leq C^{\prime}N^{-\frac{b}{2}}.$
@@
    For $z_{c}(y)=c\pm i\frac{\sqrt{y}+E_{N}^{\theta}(y)}{2}$, for any $b\in(0,1)$ and $\epsilon\in(0,1/2)$, there exists $C>0$ and $N_{0}\in\mathbb{Z}_{\geq 1}$ such that for all $N>N_{0}$ and $y\in\hat{S}^{(N),c}=[0,2N]$,
    $|\textsc{Error}_{1}^{\pm}[\kappa,z_{c}(y)]|\leq C\left(N^{-1/2}(1+\sqrt{y})^{2}+N^{-b/2}(1+\sqrt{y})^{1+2b+\epsilon}\right).$
@@
    $|\log{[\Gamma(z^{\prime}_{c}(y))]}-\log{[\Gamma(z_{c}(y))]}|\leq CN^{-\chi}(1+\sqrt{y})^{\eta}.$
@@
(3.5) $|E_{N}^{\mathcal{A}}(y)|,|E_{N}^{\mathcal{E}}(y)|\leq CN^{-\chi}(1+\sqrt{y})^{\eta}$
@@
Type (1) $=E_{N}^{+}\left[1,u+v,w+r\right]+E_{N}^{-}\left[v+t+w,v+r,u+w,u+r-t\right]-E_{N}^{+}\left[u+v+w+r\right]$
Type (2) $=E_{N}^{+}\left[\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)},\frac{-i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)}\right]$
$-E_{N}^{+}\left[v+\frac{t}{2}+\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)},v+\frac{t}{2}-\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)}\right]$
$-E_{N}^{+}\left[u-\frac{t}{2}+\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)},u-\frac{t}{2}-\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)}\right]$
$-E_{N}^{-}\left[w+\frac{t}{2}+\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)},w+\frac{t}{2}-\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)}\right]$
$-E_{N}^{-}\left[r-\frac{t}{2}+\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)},r-\frac{t}{2}-\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)}\right]$
@@
Turning our attention to $E^{\mathcal{A}}_{N}(y)$ (recall $E^{\mathcal{A}}_{N}(y)=\mathcal{A}_{N}(\sqrt{y}+E_{N}^{\theta}(y))-\mathcal{A}_{N}(\sqrt{y})$), and using the expressions for $\mathcal{A}^{\pm}[z_{1},..,z_{k}]$ from (2.13), we see that the subtraction in the definition of $E_{N}^{\mathcal{A}}$ causes all the terms involving a factor of $\pi^{2}$ to cancel. The terms which remain take one of three forms:
@@
    Terms associated to the $\log{\left[\frac{\Gamma(z)}{\sqrt{2\pi}}\right]}$ term in $\mathcal{A}^{+}:$
    The contributions from terms of this kind after cancellation take the form $\log{(\Gamma(z^{\prime}_{c}(y)))}-\log{(\Gamma(z_{c}(y)))}$ with
@@
(3.6) $N^{u+v}\sqrt{1-\frac{y}{4N}}\hat{\pi}_{t;v,w,u,r}^{(N),c}(y)=\mathfrak{p}_{t}^{c}(y)\cdot e^{\textsc{Error}_{t}^{(N),c}(y)},$
@@
$\lim_{N\to\infty}N^{u+v}\hat{\pi}_{t;w,v,u,r}^{(N),c}(y)=C_{b,d,u,v}\cdot\mathfrak{p}_{t}^{c}(y).$
@@
Assume $t\in(-2w,2r)$. The location of the atoms of $\hat{\pi}_{t}^{(N),d}(\cdot)$ associated to $\chi\in\{q^{v+\frac{t}{2}},q^{u-\frac{t}{2}}\}$ converge to those of $\mathfrak{p}_{t}^{d}(\cdot)$. When $v+\frac{t}{2}<0$ the atoms of $\hat{\pi}_{t}^{(N),d}(\cdot)$ are at $\hat{y}_{j}^{(N),v}(t)$ for $j\in[0,...,\lfloor-v-\frac{t}{2}\rfloor]$ and
$\lim_{N\to\infty}\hat{y}_{j}^{(N),v}(t)=x_{j}^{v}(t):=-4\left(v+j+\frac{t}{2}\right)^{2}.$
Similarly, when $u-\frac{t}{2}<0$ the atoms of $\hat{\pi}_{t}^{(N),d}(\cdot)$ are at $\hat{y}_{j}^{(N),u}(t)$ for $j\in[0,...,\lfloor u-\frac{t}{2}\rfloor]$ and
$\lim_{N\to\infty}\hat{y}_{j}^{(N),u}(t)=x_{j}^{u}(t):=-4\left(u+j+\frac{t}{2}\right)^{2}.$
@@
We recall the expression for $\hat{\pi}_{t}^{(N),d}(\hat{y}_{0}^{(N),v}(t))$,
$\hat{\pi}_{t}^{(N),d}(\hat{y}_{0}^{(N),v}(t))$ $=\frac{(q^{-2v-t},-q^{w+u},q^{w+r},-q^{r+u-t})_{\infty}}{(-q^{w-v},q^{u-v-t},-q^{r-v-t},q^{v+u+w+r})_{\infty}}.$
@@
$\log{\left(N^{u+v}\hat{\pi}_{t}^{(N),d}(\hat{y}_{0}^{(N),v}(t))\right)}$ $=(u+v)\log{(N)}+\mathcal{A}^{+}_{N}[-2v-t,w+r]+\mathcal{A}^{-}_{N}[w+u,r+u-t]$
$-\mathcal{A}^{+}_{N}[u-v-t,u+v+w+r]-\mathcal{A}^{-}_{N}[w-v,r-v-t]$
$+E^{+}_{N}[-2v-t,w+r]+E^{-}_{N}[w+u,r+u]$
$-E^{+}_{N}[u-v-t,u+v+w+r]-E^{-}_{N}[w-u,r-v-t].$
@@
(3.7) $N^{u+v}\hat{\pi}_{t}^{(N),d}(\hat{y}_{0}^{(N),v}(t))=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{d}(x_{0}^{v}(t))\cdot e^{\textsc{Error}_{t}^{d}(x_{0}^{u}(t))},$
@@
The remaining atoms have weights $\hat{\pi}_{t}^{(N),d}(\hat{y}_{j}^{(N),v}(t))$,
$\hat{\pi}_{t}^{(N),d}(\hat{y}_{j}^{(N),v}(t))$ $=\hat{\pi}_{t}^{(N),d}(\hat{y}_{0}^{(N),v}(t))$
@@
$\lim_{N\to\infty}N^{u+v}\hat{\pi}_{t}^{v}(y_{j}^{(N),d})$ $=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{d}(x_{j}^{v}(t))$
$\lim_{N\to\infty}N^{u+v}\hat{\pi}_{t}^{u}(y_{j}^{(N),d})$ $=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{d}(x_{j}^{u}(t))$
@@
We begin by recalling the expression for $\hat{\pi}_{s,t;v,w,u,r}^{(N),c,c}(m,r)$:
(3.9) $\hat{\pi}_{s,t;v,w,u,r}^{(N),c,c}(m,r)=\frac{1}{2N}\pi_{q^{t},q^{s}}^{(N),c,c}\left(1-\frac{r}{2N},1-\frac{m}{2N}\right)\cdot\frac{\hat{\pi}_{t;v,w,u,r}^{(N),c}(r)}{\hat{\pi}_{s;v,w,u,r}^{(N),c}(m)}$
@@
$\sqrt{1-\frac{m}{4N}}\frac{1}{2N}\pi_{q^{t},q^{s}}\left(1-\frac{r}{2N},1-\frac{m}{2N}\right)=$
$\frac{\left(q,-q^{v+w+s},q^{t-s}\right)_{\infty}\bigg{|}\left(q^{v+\frac{t}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},-q^{w+\frac{t}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},q^{i\frac{\sqrt{N}\theta_{m}}{2}}\right)\bigg{|}^{2}}{4\pi N\sqrt{\frac{m}{N}}\left(-q^{v+w+t}\right)_{\infty}\bigg{|}\left(q^{v+\frac{s}{2}+i\frac{\sqrt{N}\theta_{m}}{2}},-q^{w+\frac{s}{2}+i\frac{\sqrt{N}\theta_{m}}{2}},q^{\frac{t-s}{2}+i\frac{\sqrt{N}(\theta_{r}+\theta_{m})}{2}},q^{\frac{t-s}{2}+i\frac{\sqrt{N}(-\theta_{r}+\theta_{m})}{2}}\right)_{\infty}\bigg{|}^{2}}.$
@@
(3.10) $\log{\left(\sqrt{1-\frac{m}{4N}}\frac{1}{2N}\pi_{q^{t},q^{s}}^{(N),c,c}\left(1-\frac{r}{2N},1-\frac{m}{2N}\right)\right)}=$
$-\log{\left(4\pi N\sqrt{\frac{m}{N}}\right)}+\mathcal{A}_{N}(\sqrt{m},\sqrt{r})+E_{N}^{\mathcal{A}}(m,r)+E_{N}^{\mathcal{E}}(m,r)$
@@
$\mathcal{A}_{N}(\sqrt{m},\sqrt{r})$ $:=\mathcal{A}_{N}^{+}\left[1,t-s,v+\frac{t}{2}+i\frac{\sqrt{r}}{2},v+\frac{t}{2}-i\frac{\sqrt{r}}{2},i\sqrt{m},-i\sqrt{m}\right]$
$+\mathcal{A}_{N}^{-}\left[v+w+s,w+\frac{t}{2}+i\frac{\sqrt{r}}{2},w+\frac{t}{2}-i\frac{\sqrt{r}}{2}\right]$
$-\mathcal{A}^{+}_{N}\left[v+\frac{s}{2}+i\frac{\sqrt{m}}{2},v+\frac{s}{2}-i\frac{\sqrt{m}}{2},\frac{t-s}{2}+i\frac{\sqrt{r}+\sqrt{m}}{2}\right]$
$-\mathcal{A}^{+}_{N}\left[\frac{t-s}{2}-i\frac{\sqrt{r}+\sqrt{m}}{2},\frac{t-s}{2}+i\frac{\sqrt{r}-\sqrt{m}}{2},\frac{t-s}{2}-i\frac{\sqrt{r}-\sqrt{m}}{2}\right]$
$-\mathcal{A}^{-}_{N}\left[v+w+t,w+\frac{s}{2}+i\frac{\sqrt{m}}{2},w+\frac{s}{2}-i\frac{\sqrt{m}}{2}\right],$

### v2

@@ 3.1. Convergence of Marginal Distributions
We begin by reminding ourselves of the expression for $\widehat{\pi}_{t;v,w,u,r}^{(N),c}(\cdot)$. As before, we set $q=\exp{\left(-\frac{2}{\sqrt{N}}\right)}$. We will deal with the expression for the rescaled, time-reversed process $\widehat{\mathbb{Y}}_{t}$, and the corresponding marginal measures $\widehat{\pi}_{t;v,w,u,r}^{(N),c}(\cdot)$.
$\widehat{\pi}_{t;v,w,u,r}^{(N),c}(y)$ $=\frac{1}{2N}\pi_{q^{t}}^{(N),c}\left(1-\frac{y}{2N}\right)$ (3.3)
@@
×$\times\frac{(q,-q^{v+w+t},q^{u+v},-q^{v+r},-q^{u+w},q^{w+r},-q^{u+r-t})_{\infty}|(q^{i\sqrt{N}\theta})_{\infty}|^{2}}{|(q^{v+\frac{t}{2}+i\frac{\sqrt{N}\theta}{2}},-q^{w+\frac{t}{2}+i\frac{\sqrt{N}\theta}{2}},q^{u-\frac{t}{2}+i\frac{\sqrt{N}\theta}{2}},-q^{r-\frac{t}{2}+i\frac{\sqrt{N}\theta}{2}})_{\infty}|^{2}}.$
We will only prove a bound in the interval $y\in[0,2N]$. The way that this bound can be applied to prove the full Lemma 3.1 is by using the symmetry (2.5) to turn this bound into the proof by first proving our desired bound for $\widehat{\pi}^{(N),c}_{t;v,w,u,r}(y)$ on the interval $y\in[0,2N]$, and then noting that the desired bound on the interval $[2N,4N]$ is significantly weaker than the bound on the interval $[0,2N]$ which arises from $\widehat{\pi}^{(N),c}_{t;w,v,r,u}(4N-y)$. Therefore, we first prove the appropriate bound on (3.3) in the interval $r\in[0,2N]$, and then use this identity to reach the remaining $r\in[2N,4N].$ We take the logarithm of the marginal measure,
$\log{\left(\sqrt{1-\frac{y}{4N}}\widehat{\pi}_{t;v,w,u,r}^{(N),c}(y)\right)}$ $=-\log{\left(4\pi N\sqrt{\frac{y}{N}}\right)}+\mathcal{A}_{N}(\sqrt{y})+E^{\mathcal{A}}_{N}(y)+E^{\mathcal{E}}_{N}(y),$ (3.4)
@@
$\mathcal{A}_{N}(\sqrt{y})$ $:=\mathcal{A}^{+}_{N}\left(1,u+v,w+r,i\frac{\sqrt{y}}{2},-i\frac{\sqrt{y}}{2}\right)$
$+\mathcal{A}^{-}_{N}\left(v+w+t,v+r,u+w,u+r-t\right)$
$-\mathcal{A}_{N}^{+}\left(u+v+w+r,v+i\frac{\sqrt{y}}{2}+\frac{t}{2},u+i\frac{\sqrt{y}}{2}-\frac{t}{2},v-i\frac{\sqrt{y}}{2}+\frac{t}{2},u-i\frac{\sqrt{y}}{2}-\frac{t}{2}\right)$
$-\mathcal{A}_{N}^{-}\left(w+i\frac{\sqrt{y}}{2}+\frac{t}{2},r+i\frac{\sqrt{y}}{2}-\frac{t}{2},w-i\frac{\sqrt{y}}{2}+\frac{t}{2},r-i\frac{\sqrt{y}}{2}-\frac{t}{2}\right),$
@@
$E_{N}^{\theta}(y)$ $:=\sqrt{N}\arccos{\left(1-\frac{y}{2N}\right)}-\sqrt{y},$
are a concise way of containing the logarithm of the $q$-Pochhammer symbols in $\widehat{\pi}_{t;v,w,u,r}^{(N),c}(y)$. The notation $\mathcal{A}^{\pm}$ was defined in (2.13), and $\mathcal{A}^{\pm}_{N}$ defined in (2.15), and $E_{N}^{\theta}(y)$ arises as a concise expression for the error terms in the Taylor expansion $\sqrt{N}\theta_{y}^{(N)}=\sqrt{y}+E_{N}^{\theta}(y)$.
We can simplify the first two terms on the right-hand side of (3.4) considerably by expanding $\mathcal{A}(\sqrt{y})$ through the definitions of $\mathcal{A}^{+}_{N}(z_{1},...,z_{k})$ and $\mathcal{A}^{-}_{N}(z_{1},...,z_{k})$ from (2.15), and then canceling where appropriate. The result of that calculation yields
@@
$\log\left(N^{u+v}\sqrt{1-\frac{y}{4N}}\widehat{\pi}_{t;v,w,u,r}^{(N),c}(y)\right)$ $=\log{(\mathfrak{p}_{t}^{c}(y))}+\log{(C_{u,v,w,r})}+E_{N}^{\mathcal{A}}(y)+E_{N}^{\mathcal{E}}(y).$
@@
    For all $\eta>1$ there exists $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi>0$ such that for all $N>N_{0}$ and $y\in\widehat{S}^{(N),c}=[0,2N]$,
@@
    $|e_{1}^{\pm}(\kappa,z)|\leq C(\kappa+\kappa^{b})\leq C^{\prime}N^{-\frac{b}{2}}.$
@@
    For $z_{c}(y)=c\pm i\frac{\sqrt{y}+E_{N}^{\theta}(y)}{2}$, for any $b\in(0,1)$ and $\epsilon\in(0,1/2)$, there exists $C>0$ and $N_{0}\in\mathbb{Z}_{\geq 1}$ such that for all $N>N_{0}$ and $y\in\widehat{S}^{(N),c}=[0,2N]$,
    $|e_{1}^{\pm}(\kappa,z_{c}(y))|\leq C\left(N^{-1/2}(1+\sqrt{y})^{2}+N^{-b/2}(1+\sqrt{y})^{1+2b+\epsilon}\right).$
@@
    $|\log{\left(\Gamma(z^{\prime}_{c}(y))\right)}-\log{\left(\Gamma(z_{c}(y))\right)}|\leq CN^{-\chi}(1+\sqrt{y})^{\eta}.$
@@
$|E_{N}^{\mathcal{A}}(y)|,|E_{N}^{\mathcal{E}}(y)|\leq CN^{-\chi}(1+\sqrt{y})^{\eta}.$ (3.5)
@@
Type (1) $=E_{N}^{+}\left(1,u+v,w+r\right)+E_{N}^{-}\left(v+t+w,v+r,u+w,u+r-t\right)-E_{N}^{+}\left(u+v+w+r\right),$
Type (2) $=E_{N}^{+}\left(\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)},\frac{-i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)}\right)$
$-E_{N}^{+}\left(v+\frac{t}{2}+\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)},v+\frac{t}{2}-\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)}\right)$
$-E_{N}^{+}\left(u-\frac{t}{2}+\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)},u-\frac{t}{2}-\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)}\right)$
$-E_{N}^{-}\left(w+\frac{t}{2}+\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)},w+\frac{t}{2}-\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)}\right)$
$-E_{N}^{-}\left(r-\frac{t}{2}+\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)},r-\frac{t}{2}-\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)}\right).$
@@
Turning our attention to $E^{\mathcal{A}}_{N}(y)$ (recall $E^{\mathcal{A}}_{N}(y)=\mathcal{A}_{N}(\sqrt{y}+E_{N}^{\theta}(y))-\mathcal{A}_{N}(\sqrt{y})$), and using the expressions for $\mathcal{A}^{\pm}(z_{1},..,z_{k})$ from (2.13), we see that the subtraction in the definition of $E_{N}^{\mathcal{A}}$ causes all the terms involving a factor of $\pi^{2}$ to cancel. The terms which remain take one of three forms:
@@
    Terms associated to the $\log{\left(\frac{\Gamma(z)}{\sqrt{2\pi}}\right)}$ term in $\mathcal{A}^{+}:$
    Contributions from these terms after cancellation take the form $\log{(\Gamma(z^{\prime}_{c}(y)))}-\log{(\Gamma(z_{c}(y)))}$ with
@@
$N^{u+v}\sqrt{1-\frac{y}{4N}}\widehat{\pi}_{t;v,w,u,r}^{(N),c}(y)=\mathfrak{p}_{t}^{c}(y)\cdot e^{\textsc{Error}_{t}^{(N),c}(y)},$ (3.6)
@@
$\lim_{N\to\infty}N^{u+v}\widehat{\pi}_{t;w,v,u,r}^{(N),c}(y)=C_{b,d,u,v}\cdot\mathfrak{p}_{t}^{c}(y).$
@@
Assume $t\in(-2w,2r)$. The location of the atoms of $\widehat{\pi}_{t}^{(N),d}(\cdot)$ associated to $\chi\in\{q^{v+\frac{t}{2}},q^{u-\frac{t}{2}}\}$ converge to those of $\mathfrak{p}_{t}^{d}(\cdot)$. When $v+\frac{t}{2}<0$ the atoms of $\widehat{\pi}_{t}^{(N),d}(\cdot)$ are at $\widehat{y}_{j}^{(N),v}(t)$ for $j\in[0,...,\lfloor-v-\frac{t}{2}\rfloor]$ and
$\lim_{N\to\infty}\widehat{y}_{j}^{(N),v}(t)=x_{j}^{v}(t):=-4\left(v+j+\frac{t}{2}\right)^{2}.$
Similarly, when $u-\frac{t}{2}<0$ the atoms of $\widehat{\pi}_{t}^{(N),d}(\cdot)$ are at $\widehat{y}_{j}^{(N),u}(t)$ for $j\in[0,...,\lfloor u-\frac{t}{2}\rfloor]$ and
$\lim_{N\to\infty}\widehat{y}_{j}^{(N),u}(t)=x_{j}^{u}(t):=-4\left(u+j+\frac{t}{2}\right)^{2}.$
@@
We recall the expression for $\widehat{\pi}_{t}^{(N),d}(\widehat{y}_{0}^{(N),v}(t))$,
$\widehat{\pi}_{t}^{(N),d}(\widehat{y}_{0}^{(N),v}(t))$ $=\frac{(q^{-2v-t},-q^{w+u},q^{w+r},-q^{r+u-t})_{\infty}}{(-q^{w-v},q^{u-v-t},-q^{r-v-t},q^{v+u+w+r})_{\infty}}.$
@@
$\log{\left(N^{u+v}\widehat{\pi}_{t}^{(N),d}(\widehat{y}_{0}^{(N),v}(t))\right)}$ $=(u+v)\log{(N)}+\mathcal{A}^{+}_{N}(-2v-t,w+r),+\mathcal{A}^{-}_{N}(w+u,r+u-t)$
$-\mathcal{A}^{+}_{N}(u-v-t,u+v+w+r)-\mathcal{A}^{-}_{N}(w-v,r-v-t)$
$+E^{+}_{N}(-2v-t,w+r)+E^{-}_{N}(w+u,r+u)$
$-E^{+}_{N}(u-v-t,u+v+w+r)-E^{-}_{N}(w-u,r-v-t).$
@@
$N^{u+v}\widehat{\pi}_{t}^{(N),d}(\widehat{y}_{0}^{(N),v}(t))=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{d}(x_{0}^{v}(t))\cdot e^{\textsc{Error}_{t}^{d}(x_{0}^{u}(t))},$ (3.7)
@@
The remaining atoms have weights $\widehat{\pi}_{t}^{(N),d}(\widehat{y}_{j}^{(N),v}(t))$,
$\widehat{\pi}_{t}^{(N),d}(\widehat{y}_{j}^{(N),v}(t))$ $=\widehat{\pi}_{t}^{(N),d}(\widehat{y}_{0}^{(N),v}(t))$
@@
$\lim_{N\to\infty}N^{u+v}\widehat{\pi}_{t}^{v}(y_{j}^{(N),d})$ $=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{d}(x_{j}^{v}(t)),$
$\lim_{N\to\infty}N^{u+v}\widehat{\pi}_{t}^{u}(y_{j}^{(N),d})$ $=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{d}(x_{j}^{u}(t)),$
@@
We begin by recalling the expression for $\widehat{\pi}_{s,t;v,w,u,r}^{(N),c,c}(m,r)$:
$\widehat{\pi}_{s,t;v,w,u,r}^{(N),c,c}(m,r)=\frac{1}{2N}\pi_{q^{t},q^{s}}^{(N),c,c}\left(1-\frac{r}{2N},1-\frac{m}{2N}\right)\cdot\frac{\widehat{\pi}_{t;v,w,u,r}^{(N),c}(r)}{\widehat{\pi}_{s;v,w,u,r}^{(N),c}(m)}.$ (3.9)
@@
$\sqrt{1-\frac{m}{4N}}\frac{1}{2N}\pi_{q^{t},q^{s}}\left(1-\frac{r}{2N},1-\frac{m}{2N}\right)$
$=\frac{\left(q,-q^{v+w+s},q^{t-s}\right)_{\infty}\bigg{|}\left(q^{v+\frac{t}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},-q^{w+\frac{t}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},q^{i\frac{\sqrt{N}\theta_{m}}{2}}\right)\bigg{|}^{2}}{4\pi N\sqrt{\frac{m}{N}}\left(-q^{v+w+t}\right)_{\infty}\bigg{|}\left(q^{v+\frac{s}{2}+i\frac{\sqrt{N}\theta_{m}}{2}},-q^{w+\frac{s}{2}+i\frac{\sqrt{N}\theta_{m}}{2}},q^{\frac{t-s}{2}+i\frac{\sqrt{N}(\theta_{r}+\theta_{m})}{2}},q^{\frac{t-s}{2}+i\frac{\sqrt{N}(-\theta_{r}+\theta_{m})}{2}}\right)_{\infty}\bigg{|}^{2}}.$
@@
$\log{\left(\sqrt{1-\frac{m}{4N}}\frac{1}{2N}\pi_{q^{t},q^{s}}^{(N),c,c}\left(1-\frac{r}{2N},1-\frac{m}{2N}\right)\right)}$ (3.10)
$=-\log{\left(4\pi N\sqrt{\frac{m}{N}}\right)}+\mathcal{A}_{N}(\sqrt{m},\sqrt{r})+E_{N}^{\mathcal{A}}(m,r)+E_{N}^{\mathcal{E}}(m,r).$
@@
$\mathcal{A}_{N}(\sqrt{m},\sqrt{r})$ $:=\mathcal{A}_{N}^{+}\left(1,t-s,v+\frac{t}{2}+i\frac{\sqrt{r}}{2},v+\frac{t}{2}-i\frac{\sqrt{r}}{2},i\sqrt{m},-i\sqrt{m}\right)$
$+\mathcal{A}_{N}^{-}\left(v+w+s,w+\frac{t}{2}+i\frac{\sqrt{r}}{2},w+\frac{t}{2}-i\frac{\sqrt{r}}{2}\right)$
$-\mathcal{A}^{+}_{N}\left(v+\frac{s}{2}+i\frac{\sqrt{m}}{2},v+\frac{s}{2}-i\frac{\sqrt{m}}{2},\frac{t-s}{2}+i\frac{\sqrt{r}+\sqrt{m}}{2}\right)$
$-\mathcal{A}^{+}_{N}\left(\frac{t-s}{2}-i\frac{\sqrt{r}+\sqrt{m}}{2},\frac{t-s}{2}+i\frac{\sqrt{r}-\sqrt{m}}{2},\frac{t-s}{2}-i\frac{\sqrt{r}-\sqrt{m}}{2}\right)$
$-\mathcal{A}^{-}_{N}\left(v+w+t,w+\frac{s}{2}+i\frac{\sqrt{m}}{2},w+\frac{s}{2}-i\frac{\sqrt{m}}{2}\right),$